The density of mercury is 13.6grams per centimetres cubed. Find the volume of 2720grams of mercury in centimetres cubed and metres cubed
1
Expert's answer
2011-07-06T07:40:25-0400
The volume of the mercury can be calculated be the formula
V = m/ ρ
V = 2720 [g] / 13.6 g/cm3 = 200 cm3
In one cubic meter there are 106 cm3, therefore 200 cm3 = 200/106 m3 = 2 x 10 -4 m3.